OpenStudy (anonymous):

Please help me out. :( A baseball is thrown vertically upward with a speed, as it leaves the hand, of 30 m/s. a) how long will it rise? b) how high will it rise? c) how long after it leaves the hand will it return to the starting point? d) when will its speed be 20 m/s?

OpenStudy (rational):

\[\large v(t) - v(0) = \int \limits_0^t a~ dt\]

OpenStudy (anonymous):

yes.. slight.

OpenStudy (rational):

you know \(\large v(0) = 30\) and \(a\) is constant can you integrate ?

OpenStudy (rational):

\[\large v(t) - 30 = at\]

OpenStudy (anonymous):

am, we don't use integration for solving this..

OpenStudy (rational):

Okay, you want to solve it using kinematics equations is it ?

OpenStudy (anonymous):

yes. :) you got it. :)

OpenStudy (rational):

\[\large v = u + at\]

OpenStudy (rational):

you know \(u = 30\) and at highest point \(v = 0\) plug them and solve \(t\)

OpenStudy (rational):

that gives you the time for which it rises

OpenStudy (anonymous):

what is "u" in the equation v = u + at ?

OpenStudy (rational):

u : initial velocity v : final velocity your notes uses different letters is it ?

OpenStudy (anonymous):

ahh.. got it. :) ^_^ then, can i plug it now?

OpenStudy (rational):

yes

OpenStudy (anonymous):

okay.. i got 3.058 s for letter a. :)

OpenStudy (rational):

what value are you using for g ?

OpenStudy (rational):

9.8 or 10 ?

OpenStudy (anonymous):

i use.... -9.81

OpenStudy (rational):

g = 9.81 a = -g = -9.81

OpenStudy (rational):

g is a positive constant always.

OpenStudy (anonymous):

what? it's direction is upward right?

OpenStudy (rational):

g is a constant whose value is 9.81 or whatever it is, its always positive. a = -g

OpenStudy (rational):

since it is acting down

OpenStudy (anonymous):

ahh.. got it.. :) is my answer correct?

OpenStudy (rational):

Correct! for part b, you may use below : Maximum height = \(\dfrac{u^2}{2g}\)

OpenStudy (rational):

just plugin the numbers and evaluate

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

why ?....you have the values give the machine(formula) its input and it will feed you back with yar desired output

OpenStudy (anonymous):

please explain it to me.... why did he use that formula... thankyou.

OpenStudy (anwaarullah):

nitial velocity (u) = 30 velocity at the highest point = 0 ( becayse velocity decreases when something is thrown up and becomes 0 at the maximam point ) acceleration = -9.81 / 9.81 downwards (a) V= u+at 0 = 30+ (-9.81)*t t =3.05s (b) V^2=u^2+2as 0 = 30^2 +2* (-9.81) * s s = 45.9m (c) t*2 3.05*2 =6.1 s (d) V=u +at 20= 30= (-9.81)*t t = 1.01s methods are correct,,but smtimes there might be mistakes in da calculations because i did it in a hurry ! anyway this is the method to do the qstion :) hope it helped

OpenStudy (rational):

you don't like integral you don't like formula what do you want ? :o

OpenStudy (anonymous):

have you read about the formula's of the motion if yes ...he used some simple formulas and replaced the variable of acceleration with g .....and got the resulting formula for the hieght ..DO you want to see how it is derived? @itsmeeQUENNIE

OpenStudy (anonymous):

no.. it's not what you think @rational .. i just dont get it. sorrry. :(

OpenStudy (anonymous):

Do you get this thing posted by @anwaarullah : (b) V^2=u^2+2as 0 = 30^2 +2* (-9.81) * s s = 45.9m in the above formula just make few changes for the acceleration parameter and you can simple derive the the formula for max. height

OpenStudy (anwaarullah):

you have to write down coupl of time the answer and formulas,,,without lookng at the answer,,,then your mind will strat pickng it,,,i must say

OpenStudy (anonymous):

@letterb .. i don't get it..

OpenStudy (anonymous):

not direct explianation but gives you some direct hints

OpenStudy (anonymous):

ahh.. got it.. thanks @Adjax

OpenStudy (anwaarullah):

letter c,,,,its double distance from downward to upward and than donward,,,, thts why multiply with 2* 3.5 i guess in part ,(a)

OpenStudy (anonymous):

ahhh.. :))) got it.. then in letter b... 20 m/s will be the final velocity, right?

OpenStudy (anonymous):

letter d rather...

OpenStudy (anwaarullah):

yup you are right,,b letter,,d asking about time??? now we have initial and final velocities,,,just put them in formuls,,,,but remeber you have to learn,,,memeorise the formula whtever you do,,,,note down,,,writ down again and again,,,and get the theme of question,,,,, got it,,,,,but writing down is must

OpenStudy (anonymous):

yes.. thank you. i'll do it. :) thank you very much.

OpenStudy (anwaarullah):

no problem lol,,,wlecom any time but when available,,,remeber writing down question,,,special math or physcs nomericls,,,,are very helpful to write down,,,,,,got it

OpenStudy (anonymous):

aye aye captain. :)

OpenStudy (anwaarullah):

captain,,,,huh,,,,when i used to study,,,i write down alll formulas,,,,,whther math or phy

OpenStudy (anonymous):

yeah right.. but wait.. when you will use that formulas? how will you know that you will use that formulas?

OpenStudy (anonymous):

that;s fairly easy you need your intuition..JUst know what variable you currently have at the hand and user the simplex formula on whioch plug-chuggin will get you the answer

OpenStudy (anwaarullah):

from v^= U^2 +2as you cant calculate time???could you you.... similarly likwise

OpenStudy (anonymous):

ahh.. okay.. when you will use.... v = u + at ?

OpenStudy (anwaarullah):

let suppose thy are asking about time,,,like in part (a) ,,,you could see

OpenStudy (anwaarullah):

similary you cn nt count distance from abov formula which you mention recently

OpenStudy (anonymous):

ahh.. got it.. but, how about s = ut + 1/2 at^2, when you will use that?

OpenStudy (anwaarullah):

wher is (g) and (a )are acceleration if you know!!!! g when height ect mention,,,,,a,,,when horizntal or ground mention